OpenStudy (anonymous):

evaluate the integral e^x sin(1+e^x)dx

OpenStudy (anonymous):

∫ (e^x) sin(e^x) dx = note that you have both the sine argument and its derivative, thus let e^x = u then, differentiate both sides, yielding: d(e^x) = du → (e^x) dx = du then, substituting, you get: ∫ (e^x) sin(e^x) dx = ∫ sin(e^x) (e^x) dx = ∫ sin(u) du = - cos(u) + c finally, substituting back u = (e^x): ∫ (e^x) sin(e^x) dx = - cos(e^x) + c

OpenStudy (anonymous):

please vote for my design http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

@TheAni-Trix please vote

OpenStudy (anonymous):

oh okay thank you, i tired the problem myself and got the same answer so thanks ^_^

OpenStudy (anonymous):

please please please

OpenStudy (anonymous):

@TheAni-Trix